OpenStudy (anonymous):

Given normal random variable Z , with mean μ=17.1 and standard deviation σ=6.2 , we obtain a sample of Z of size 60 . Consider the new random variable Xˉ which is the average of our sample. Find the interval [μXˉ−a,μXˉ+a] symmetric about μXˉ satisfying P(Xˉ∈[μXˉ−a,μXˉ+a])=0.95

OpenStudy (anonymous):

what's the distribution of the sample mean for a normal r.v.?

OpenStudy (anonymous):

i dont for sure

OpenStudy (anonymous):

\[X^{-} \sim N \left( \mu,\frac{ \sigma^2 }{n } \right)\]

OpenStudy (anonymous):

meam is 17.1?

OpenStudy (anonymous):

is \(X^-\) supposed to be \(\bar{X}\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\bar{X} in latex

OpenStudy (anonymous):

\overline{X} may work better, let's see \(\overline{X}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{ \overline{X}-\mu }{ \Large\frac{ \sigma }{ \sqrt{n} } }\]is a standard normal

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

mu=17.1, sigma = 6.2 and n=60 does that men x= .95?

OpenStudy (anonymous):

\[\mu \overline{X}+a=18.6688...\]

OpenStudy (anonymous):

\[\mu \overline{X} - a = 15.5311...\]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

i think your \(\mu\overline{X}\) is really \(\mu_{\overline{X}}\)